# supplementary

### v1

@@ 1. Introduction
We exploit conjugate distribution theory for hierarchical models to account for the differences between the values of the spatial effects as well as the posterior probabilities of such differences exceeding a certain threshold. In this regard, we extend the approaches in lu2005bayesian and fitzpatrickEcologicalBoundaryDetection2010, who use only posterior means or spatial residuals to detect differences, to full Bayesian inference using posterior distributions to quantify uncertainty. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while accounting for false discoveries with the Bayesian FDR (mullerOptimalSampleSize2004).
@@ 2.2. Random Effects under an Augmented Linear System
If $\rho\in(0,1)$ is fixed, then we can draw exact samples from the joint posterior. Denote $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$ and $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}$. Familiar calculations yield $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)\left(A^{-1}+A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}XA^{-1}\right)$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$. Similarly, from $\mathbb{E}[\gamma_{\star}\,|\,y,\sigma^{2},\rho]=M_{\star}m_{\star}$ we obtain $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=A^{-1}(X^{\mathrm{\scriptscriptstyle T}}y+(1-\rho)m_{0})-A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}(y-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}y-(1-\rho)XA^{-1}m_{0})$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(y-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}y-(1-\rho)XA^{-1}m_{0})$. Since $\gamma=\sigma\sqrt{\rho}\phi$, we have
@@
A flat prior on $\beta$ considerably simplifies $\mathbb{E}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$ and $\operatorname{Var}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$. If $\pi(\gamma,\beta,\sigma^{2})=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})$, then $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-H$, where $H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$, $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}y-(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}(I-H)y$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}+\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I-H)y$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$. Since $\gamma=\sigma\sqrt{\rho}\phi$, we obtain

### v2

@@ 1. Introduction
We exploit conjugate distribution theory for hierarchical models to account for the differences between the values of the spatial effects as well as the posterior probabilities of such differences exceeding a certain threshold. In this regard, we extend the approaches in lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to full Bayesian inference using posterior distributions to quantify uncertainty. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while accounting for false discoveries with the Bayesian FDR (mullerOptimalSampleSize2004).
@@ 2.2. Random Effects under an Augmented Linear System
If $\rho\in(0,1)$ is fixed, then we can draw exact samples from the joint posterior. Denote $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$ and $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}$. Familiar calculations yield $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)\left(A^{-1}+A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}XA^{-1}\right)$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$. Similarly, from $\mathbb{E}[\gamma_{\star}\,|\,y,\sigma^{2},\rho]=M_{\star}m_{\star}$ we obtain $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(1-\rho)/\rho\cdot A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}V_{\phi}^{-1}y+(1-\rho)A^{-1}\left(X^{\mathrm{\scriptscriptstyle T}}B^{-1}X+I_{n}\right)m_{0}$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}})y-(1-\rho)B^{-1}XA^{-1}m_{0}$. Since $\gamma=\sigma\sqrt{\rho}\phi$, we have
@@
A flat prior on $\beta$ considerably simplifies $\mathbb{E}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$ and $\operatorname{Var}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$. If $\pi(\gamma,\beta,\sigma^{2})=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})$, then $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-H$l, where $H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$, $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(1-\rho)\cdot(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}V_{\phi}^{-1}y/\rho$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}+\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I-H)y$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$. Since $\gamma=\sigma\sqrt{\rho}\phi$, we obtain